CAT Chat Minutes

July 10, 1998

Information/Follow-up Items

1. For the past month, the vending machines have not been kept up. They have had no food in them. Can something be done?

We are requesting the vendor to monitor the status of the vending machines more frequently.

2. Many of our users have been eating outside during the nice weather. Because of this, there is more food put into the trash receptacles, and the raccoons have been kept very busy. Can something be done?

This has been an on-going problem. We have requested the ANL Grounds personnel to be more proactive in emptying the receptacles.

3. Mosquitoes are getting quite numerous on the floor. Can ANL get an exterminator to come in?

We cannot fog on the floor. The chemicals used are detrimental to electronic equipment.

4. Does ANL have a mosquito abatement program as some of the suburbs do?

No, the lab does not have an abatement program.

5. During the next run, you have plans to test top-up mode. Has a schedule been determined?

No top-up runs are planned during the next user run. We plan to use Machines Studies and Shielding Verification times for performing initial top-up studies.

6. Do you have anymore information on the computer break-in?

The news that the Qpopper software was vulnerable reached APS on June 29. The patch for the software was made available at 3 AM on June 30. We were attacked at 4 AM on June 30. We discovered the problem at 8:30 AM on June 30. It was determined that at least one system program had been replaced with a fake. The systems were removed from the internet, and all system files restored from tape at 5 PM on June 30. The patched Qpopper software was installed. Additional security measures were also put in place.

The APS security measures had been tested and approved in April by ECT and the DOE OSE Site Profile inspection team. No security problems were found at that time. This was an example of a new form of attack on an established vendors product.

The APS has been probed several times in the last week. These probes have not been successful.

APS Operational Status Update

We had been running at ~ 97% up to this morning. Overall, for the run we are still at ~ 94%. We lost eight hours this morning due to an rf trip in the storage ring, which also tripped off the injector rf. Repairing the rf systems and additional problems with the storage ring main dipole power supply contributed to the largest block of downtime during this eight week run.

Tuesday July 14, 1998, work will be done on the software for the HVAC. Please do not work in your hoods from 1100 to 1300 hours on Tuesday. Thursday, the auto transfer switch will be worked on. The work will be done in two parts. One sequence will be at 0800 to 0915 and will affect sectors 34 through 13. The second sequence will be done from 1230 to 1315 and will affect sectors 30 to 17. There will be a short interruption while they switch. Work on the X6 switch gear will be done July 18 and will take ~ 12 hours. The work will only impact PNC, but B. Ferry will make arrangements to provide emergency power for sector 20.

Information

We have found that only a few users require small quantities of dry ice and only occasionally. If anyone has the need for dry ice, please contact B. Ferry at x 9495, preferably a few weeks in advance of your need.

If you want to have your computers scanned for possible problems with break-ins during the shutdown, please contact B. Ferry at x 9495. We highly recommend that everyone do this.